OpenStudy (anonymous):

An airplane flying into a headwind travels 1160 miles in 4 hours and 50 minutes. On the return flight, the distance is traveled in 4 hours. Find the airspeed of the plane and the speed of the wind, assuming that both remain constant.

OpenStudy (anonymous):

https://www.youtube.com/watch?v=6FmWAvyTYBA&feature=endscreen&NR=1 this will really help

OpenStudy (nurali):

Let S = the airplane's speed Let W = the wind speed 4 hours and 50 minutes = 4.8 hours Into the wind, (S)(4.8 hours) - (W)(4.8 hours) = 1160 miles With the wind, (S)(4.0) + (W)(4.0) = 1160 miles Using this equation first, divide both sides by 4.0, to yield: S + W = 1160/4.0 = 290 Therefore, S = 290 - W Substituting this value for S into the first equation yields: (290 - W)(4.8) - (W)(4.8) = 1160 Again, divide both sides by 4.8 to yield: 290 - W - W = 242 Rearranging gives us: 290 - 242 = 2W Therefore; W = 24 Because S = 290 - W, S = 290 - 24 = 266

OpenStudy (whpalmer4):

\[d = rt\]where \(d\) is distance, \(r\) is rate or speed, and \(t\) is time. Let the airplane's speed be \(v_a\) and the wind \(v_w\). When the airplane flies with the wind (tailwind) the speed over the ground will be \(v_a+v_w\). When the airplane flies against the wind (headwind) the speed over the ground will be \(v_a-v_w\). We know that the distance in both cases is the same, 1160 miles. Flying into the headwind: 1160 miles = 4 hours 50 minutes * \((v_a-v_w)\) Flying with the wind: 1160 miles = 4 hours * \((v_a+v_w)\) Convert the times to minutes (or decimal hours), expand the two equations, and solve.